# DAWN-FM: Data-Aware and Noise-Informed Flow Matching for Solving Inverse Problems

### v2

@@ Appendix C Experimental Settings
Optimizer Adam [Kingma and Ba, 2014]
@@
Table 4: Experimental details for the image deblurring task for Diffusion model [Chung et al., 2022a].
@@
Optimizer Adam [Kingma and Ba, 2014]
@@
Table 5: Experimental details for the image deblurring task for InverseUNetODE [Eliasof et al., 2024].
@@
Optimizer Adam [Kingma and Ba, 2014]
@@
Table 6: Experimental details for the tomography task for our methods DAW-FM and DAWN-FM.
@@
[Yang et al., 2023]. The dimension of images in both these datasets
@@
DAW-FM: UNet with 5 levels with [3, 16, 32. 64, 128] filters
@@
Optimizer Adam [Kingma and Ba, 2014]
@@
Appendix D Visualization
Figure 7: Comparison between different methods for the image deblurring task on the MNIST dataset for noise levels $p=5\%$ and 7% added to the blurred data.
Figure 8: Comparison between different methods for the image deblurring task on the STL10 dataset for noise levels $p=5\%$ and 7% added to the blurred data.
Figure 9: Comparison between different methods for the image deblurring task on the CIFAR10 dataset for noise levels $p=5\%$ and 7% added to the blurred data.
Figure 10: Comparison between DAW-FM and DAWN-FM methods for the tomography task on the OrganAMNIST dataset for noise levels $p=5\%$ and 7% added to the sinogram data.
Figure 11: Comparison between DAW-FM and DAWN-FM methods for the tomography task on the OrganCMNIST dataset for noise levels $p=5\%$ and 7% added to the sinogram data.

### v3

@@ 8. Experimental Settings
Optimizer Adam [kingma2014adam]
@@
Table 5. Experimental details for the image deblurring task for Diffusion model [chung2022diffusion].
@@
Optimizer Adam [kingma2014adam]
@@
Table 6. Experimental details for the image deblurring task for InverseUNetODE [eliasof2024completedeeplearningmethod].
@@
Optimizer Adam [kingma2014adam]
@@
Table 7. Experimental details for the tomography task for our methods DAW-FM and DAWN-FM.
@@
[medmnistv2]. The dimension of images in both these datasets
@@
DAW-FM: UNet with 5 levels with [1, 16, 32, 64, 128] filters
@@
Optimizer Adam [kingma2014adam]
@@
9. Visualization
Figure 8. Comparison between different methods for the image deblurring task on the MNIST dataset for noise levels $p=5\%$ and 7% added to the blurred data.
Figure 9. Comparison between different methods for the image deblurring task on the STL10 dataset for noise levels $p=5\%$ and 7% added to the blurred data.
Figure 10. Comparison between different methods for the image deblurring task on the CIFAR10 dataset for noise levels $p=5\%$ and 7% added to the blurred data.
Figure 11. Comparison between different methods the tomography task on the OrganAMNIST dataset for noise levels $p=5\%$ and 7% added to the sinogram data.
Figure 12. Comparison between different methods the tomography task on the OrganCMNIST dataset for noise levels $p=5\%$ and 7% added to the sinogram data.